# The 3D-PC: a benchmark for visual perspective taking in humans and machines

### v1

@@ 1 Introduction
In his theory of cognitive development, Piaget posited that human children gain the ability to predict which objects are visible from another viewpoint before the age of 10 [1, 2]. This “Visual Perspective Taking” (VPT) ability is a foundational feature of human intelligence and a behavioral marker for the theory of mind [3]. VPT is also critical for safely navigating through the world and socializing with others (Fig. 1A). While VPT has been a focus of developmental psychology research since its initial description [1, 4, 5] (Fig. 1B), it has not yet been studied in machines.
@@
Figure 1: Visual Perspective Taking (VPT) is the ability to analyze scenes from different viewpoints. (A) Humans rely on VPT to anticipate the behavior of others. We expect that this ability will be essential for creating the next generation of AI assistants that can accurately anticipate human behavior (images are CC BY-NC). (B) VPT has been studied in developmental psychology since the mid-20th century using cartoon or highly synthetic stimuli. For example, Piaget’s “Three Mountains Task” asks observers to describe the scene from the perspective of a bear (image from  [6]). (C) Here, we use Gaussian Splatting [7] to develop a 3D scene generation pipeline for the 3D perception challenge (3D-PC), to systematically compare 3D perception capabilities of human and machine vision systems. (D) The 3D-PC tests 1. Object depth perception, and 2. VPT.
@@
One of the more surprising results in deep learning has been the number of concomitant similarities to human perception exhibited by deep neural networks (DNNs), trained on large-scale static image datasets [8, 9]. For example, DNNs now rival or surpass human recognition performance on object recognition and segmentation tasks [10, 11, 12], and are the state-of-the-art approach for predicting human neural and behavioral responses to images [13]. There is also a growing number of reports indicating that DNNs trained with self-supervision or for object classification learn to encode 3D properties of objects and scenes that humans are also sensitive to, such as the depth and structure of surfaces [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Are the emergent capabilities of DNNs for 3D vision sufficient for solving VPT tasks?
@@
Here, we introduce the 3D perception challenge (3D-PC) to address this question and systematically compare 3D perceptual capabilities of humans and DNNs. The 3D-PC evaluates observers on (Fig 1): 1. identifying the order of two objects in depth (depth order), 2. predicting if one of two objects can “see” the other (VPT-basic), and 3. another version of VPT that limits the effectiveness of “shortcut” solutions [24] (VPT-Strategy). The 3D-PC is distinct from existing psychological paradigms for evaluating VPT [1, 4, 5] and computer vision challenges for 3D perception [23, 20] in two ways. First, unlike small-scale psychology studies of VPT, the 3D-PC uses a novel “3D Gaussian Splatting” [7] approach which permits the generation of endless real-world stimuli. Second, unlike existing computer vision challenges, our approach for data generation means that the 3D-PC tests and counterbalances labels for multiple 3D tasks on the exact same images, which controls for potential confounds in analysis and interpretation. We expect that DNNs which rival humans on the 3D-PC will become ideal models for a variety of real-world applications where machines must anticipate human behavior in real-time, as well as for enriching our understanding of how brains work (Fig. 1A).
@@
We built the 3D-PC and used it to evaluate 3D perception for human participants and 327 different DNNs. The DNNs we tested represented each of today’s leading approaches, from Visual Transformers (ViT) [25] trained on ImageNet-21k [26] to ChatGPT4 [27] and Stable Diffusion 2.0 [28].
@@
  • Fine-tuning the zoo of DNNs on VPT-basic boosted their performance to near human level. However, the performance of the DNNs — but not humans — dropped back to chance on VPT-Strategy.
  • Our findings demonstrate that the visual strategies necessary for solving VPT do not emerge in DNNs from large-scale static image training or after directly fine-tuning on the task. We release the 3D-PC data, code, and human psychophysics at https://github.com/serre-lab/VPT to support the development of models that can perceive and reason about the 3D world like humans.
@@ 2 Related work
The visual perception of 3D properties is a fundamentally ill-posed problem [29, 30], which forces biological visual systems to rely on a variety of assumptions to decode the structure of objects and scenes. For example, variations in the lighting, texture gradients, retinal image disparity, and motion of an object all contribute to the perception of its 3D shape. 3D perception is further modulated by top-down beliefs about the structure of the world, which are either innate or shaped by prior sensory experiences, especially visual and haptic ones. In other words, humans learn about the 3D structure of the world in an embodied manner that is fundamentally different than how DNNs learn. In light of this difference, it would be remarkable if DNNs could accurately model how humans perceive their 3D world.
@@
VPT was devised to understand how capabilities for reasoning about objects in the world develop throughout the course of one’s life. At least two versions of VPT have been introduced over the years [31, 32]. The version of VPT that we study here — known in the developmental literature as “VPT-1” — is the more basic form, which is thought to rely on automatic feedforward processing in the visual system [31]. In light of the well-documented similarities between feedforward processing in humans and DNNs [13, 33], we reasoned that this version of VPT would maximize the chances of success for today’s DNNs.
3D perception in DNNs trained on static images.
As deep neural networks (DNNs) have increased in scale and training dataset size over the past decade, their performance on essentially all visual challenges has improved. Surprisingly, this “scale-up” has also led to the emergence of 3D perceptual capabilities. For example, DNNs trained with a variety of self-supervised learning techniques on static image datasets learn to represent the depth, surface normals, and 3D correspondence of features in scenes [15, 16, 17, 18, 19, 20, 21, 22, 23]. While similarities between DNNs and human 3D perception have yet to be evaluated systematically, it has been shown that there are differences in how the two reason about the 3D shape of objects [34]. The 3D-PC complements prior work by systematically evaluating which aspects of human 3D perception today’s DNNs can and cannot accurately represent.
@@
Over recent years, DNNs have grown progressively more accurate as models of human vision for object recognition tasks [10, 24]. At the same time, these models which succeed as models of human object recognition struggle to capture other aspects of visual perception [35] including contextual illusions [36], perceptual grouping [37, 38], and categorical prototypes [39]. There are also multiple reports showing that DNNs are growing less aligned with the visual strategies of humans and non-human primates as they improve on computer vision benchmarks [40, 41, 42]. The 3D-PC provides another axis upon which the field can evaluate DNNs as models of human vision.
@@ 3 Methods
To enable a fair comparison between human observers’ and DNNs’ 3D perceptual capabilities, we designed the 3D-PC framework with two goals: 1. posing different 3D tasks on the same set of stimuli, and 2. generating a large number of stimuli to properly train DNNs on these tasks. We achieved these goals by combining 3D Gaussian Splatting [7], videos from the Common Objects in 3D (Co3D) [43] dataset, and Unity [44, 45] into a flexible data-generating framework.
@@
Our approach makes it possible to generate an unlimited number of visual stimuli that test an observer’s ability to solve complementary 3D perception tasks (depth order and VPT) while keeping visual statistics constant and ground truth labels counterbalanced across tasks. For the version of 3D-PC used in our evaluation and released publicly at https://github.com/serre-lab/VPT, the depth order and VPT-basic tasks are posed on the same set of 7,480 training images of 20 objects and scenes, and a set of 94 test images of 10 separate objects and scenes (Fig. 2). We held out a randomly selected 10% of the training images for validation and model checkpoint selection.
To build the VPT-Strategy task, we rendered images where we fixed the scene camera while we moved the green camera and red ball objects to precisely change the line-of-sight between them from unobstructed to obstructed and back. We reasoned that this experiment would reveal if an observer adopts the visual strategy of taking the perspective of the green camera, which is thought to be used by humans [31], from other strategies that relied on less robust feature-based shortcuts. This dataset consisted of a test set of 100 images for 10 objects and scenes that were not included in depth order or VPT-basic.
@@
Participants were shown instructions for one of the 3D-PC tasks, then provided 20 examples to ensure that they properly understood it (Appendix Fig A.1). These examples were drawn from the DNN training set. Each experimental trial consisted of the following sequence of events overlaid onto a white background: 1. a fixation cross displayed for $1000ms$; 2. an image displayed for $3000ms$, during which time the participants were asked to render a decision. Participants pressed one of the left or right arrow keys on their keyboards to provide decisions.
@@
We evaluated a wide range of DNNs on the 3D-PC, which represented the leading approaches for object classification, self-supervised pretraining, image generation, depth prediction, and vision language modeling (VLM). Our zoo includes 317 DNNs from PyTorch Image Models (TIMM) [46], ranging from classic models like AlexNet [47] to state-of-the-art models like EVA-02 [48] (see Appendix 1 for the complete list). We added foundational vision models like MAE [49], DINO v2 [50], iBOT [51], SAM [52], and Midas [15] (obtained from the GitHub repo of [23]). We also included Depth Anything [53], a foundational model 3D scene analysis and depth prediction [23], as well as the Stable Diffusion 2.0 [28] image generation model. Finally, we added state-of-the-art large vision language models (VLMs) ChatGPT4 [27], Gemini [54], and Claude 3 [55]. We evaluated a total of 327 models on the 3D-PC.
@@
We evaluated all models except for the VLMs on the depth order and VPT-basic tasks in this challenge by training linear probes on image embeddings from their penultimate layers. Linear probes were trained using PyTorch [56] for 50 epochs, a $5e$-4 learning rate, and early stopping (see Appendix A.5 for details). Training took approximately 20 minutes per model using NVIDIA-RTX 3090s. We tested the Stable Diffusion 2.0 model by adopting the evaluation method used in [14] (see Appendix A.7 for details). We evaluated the VLMs by providing them the same instructions and training images (along with ground truth labels) given to humans, then recording their responses to images from each task via model APIs.
@@ 4 Results
Human participants were on average 74.73% accurate at determining the depth order of objects, and 86.82% accurate at solving the VPT-basic task (Fig. 3; $p<0.001$ for both; statistical testing done through randomization tests [57]). Humans were also significantly more accurate at solving VPT-basic than they were at the depth order task.
@@
DNNs showed the opposite pattern of results on depth order and VPT-basic tasks as humans after linear probing or prompting (Fig. 4): 15 of the DNNs we tested fell within the human accuracy confidence interval on the depth order task, and three even outperformed humans (Fig. 4A). In contrast, while humans were on average 86.82% accurate at VPT-basic, the DNN which performed the best on this task, the ImageNet 21K-trained beit [58], was 53.82% accurate. Even commercial VLMs struggled on VPT-basic and were around chance accuracy (ChatGPT4: 52%, Gemini: 52%, and Claude 3: 50%). The depth order task was significantly easier for DNNs than VPT-basic ($p<0.001$), which is the opposite of humans (Fig. 4B).
@@
What drives the development of 3D perception in DNNs trained on static images? We hypothesized that as DNNs scale up, they learn ancillary strategies for processing natural images, including the ability to analyze the 3D structure of scenes. To investigate this possibility, we focused on the TIMM models in our DNN zoo. These models have previously been evaluated for object classification accuracy on ImageNet, which we used as a stand-in for DNN scale [41, 42, 40]. Consistent with our hypothesis, we found a strong and significant correlation between DNN performance on ImageNet and depth order task accuracy ($\rho=0.66,p<0.001$, Fig. 4C). Despite the very low accuracy of DNNs on VPT-basic, there was also a weaker but still significant correlation between performance on this task and ImageNet ($\rho=0.34,p<0.001$, the difference in correlations between the tasks is $\rho=0.32,p<0.001$; Fig. 4D). These results suggest that monocular depth cues develop in DNNs alongside their capabilities for object classification  [^1] More work is needed to identify a causal relationship between the development of monocular depth cues and object recognition accuracy.. However, the depth cues that DNNs learn are poorly suited for VPT.
Figure 6: Even DNNs fine-tuned on VPT-basic fail on VPT-Strategy. (A) To better characterize the strategy used by humans and DNNs to solve VPT, we devised a new test, VPT-Strategy, in which the green camera and red ball are moved through a scene while holding the scene camera and a centrally-positioned object still. This task is easily solvable if an observer estimates the line-of-sight of the green camera; other strategies, such as those that rely on specific image features (feature based), may be less effective. (B) Examples of VPT-Strategy stimuli along with the ground-truth label (top-row) and predictions by a ViT large after linearly probing or fine-tuning for VPT-basic (bottom-row). Decision attribution maps from each version of the ViT large, derived from “smooth gradients” [59], are overlaid onto bottom-row images (purple/blue=linearly probed, yellow/green=fine-tuned). The fine-tuned ViT locates the green camera and red ball but renders incorrect decisions. (C) DNNs fine-tuned on VPT-basic fail to solve VPT-Strategy; they rely on a brittle feature-based strategy. Humans, on the other hand, are 87% accurate; they likely estimate line-of-sight.
@@
One possible explanation for the failure of today’s DNNs on VPT-basic is that the task requires additional cues for 3D vision that cannot be easily learned from static images. To explore this possibility, we fine-tuned each of the TIMM models in our DNN zoo to solve depth order and VPT-basic (Fig. 5A). There was still a significant difference between DNN performance on the two tasks (Fig. 5B, $p<0.001$), but fine-tuning caused 97% of the DNNs to exceed human accuracy on depth order, and four of the DNNs to reach human accuracy on VPT-basic. DNN performance on the tasks more strongly correlated with ImageNet accuracy after fine-tuning than linear probing (compare Fig. 5C/D and Fig. 4C/D). We also compared the errors these DNNs made on both tasks to humans. We found nearly all of the fine-tuned DNNs were aligned with humans on depth order, and a handful were aligned with humans on VPT-basic (Fig. A.3).
@@
The ability of DNNs to reach human-level performance on visual tasks by adopting strategies that are different from humans has been well-documented [40, 41, 42]. Thus, we devised a new experiment to understand if DNNs learn to solve VPT in the same way as humans do after fine-tuning. In developmental psychology, it has been proposed that humans estimate the line-of-sight of objects for VPT because they respond in predictable ways after the positions of objects in a scene are slightly adjusted [32, 31]. Inspired by this psychological work, we created the VPT-Strategy task to evaluate the types of visual strategies used by DNNs and humans to solve VPT (Fig. 6A).
VPT-Strategy has observers solve the VPT task on a series of images rendered from a fixed camera viewpoint as the green camera and red ball are moved incrementally from one side of the screen to the other, passing by an occluding object in the process. This means that we can precisely map out the moments at which the green camera has a clear view of the red ball, when that view is occluded, and when the view becomes unoccluded once more. DNNs behave differently than humans on this task: humans were 87% accurate, but the highest performing DNN, the Swin Transformer [60] trained on ImageNet-21k, was only 66% accurate (Fig. 6B, C). In other words, while DNNs can be fine-tuned to approach human accuracy on VPT-basic, the strategy they learn is brittle, generalizes poorly, and is likely ill-suited for reasoning about the 3D world.

### v2

@@ 1 Introduction
In his theory of cognitive development, Piaget posited that human children gain the ability to predict which objects are visible from another viewpoint before the age of 10 (Piaget et al., 1956; Frick et al., 2014). This “Visual Perspective Taking” (VPT) ability is a foundational feature of human intelligence and a behavioral marker for the theory of mind (Aichhorn et al., 2006). VPT is also critical for safely navigating through the world and socializing with others (Fig. 1A). While VPT has been a focus of developmental psychology research since its initial description (Piaget et al., 1956; Bukowski, 2018; Martin et al., 2019) (Fig. 1B), it has not yet been studied in machines.
@@
Figure 1: Visual Perspective Taking (VPT) is the ability to analyze scenes from different viewpoints. (A) Humans rely on VPT to anticipate the behavior of others. We expect that this ability will be essential for creating the next generation of AI assistants that can accurately anticipate human behavior (images are CC BY-NC). (B) VPT has been studied in developmental psychology since the mid-20th century using cartoon or highly synthetic stimuli. For example, Piaget’s “Three Mountains Task” asks observers to describe the scene from the perspective of a bear (image from  Bruce et al. (2017)). (C) Here, we use Gaussian Splatting (Kerbl et al., 2023) to develop a 3D scene generation pipeline for the 3D perception challenge (3D-PC), to systematically compare 3D perception capabilities of human and machine vision systems. (D) The 3D-PC tests 1. Object depth perception, and 2. VPT.
@@
One of the more surprising results in deep learning has been the number of concomitant similarities to human perception exhibited by deep neural networks (DNNs), trained on large-scale static image datasets (Yamins & DiCarlo, 2016; Yamins et al., 2014). For example, DNNs now rival or surpass human recognition performance on object recognition and segmentation tasks (Geirhos et al., 2021; Linsley* et al., 2021; Lee et al., 2017), and are the state-of-the-art approach for predicting human neural and behavioral responses to images (Serre, 2019). There is also a growing number of reports indicating that DNNs trained with self-supervision or for object classification learn to encode 3D properties of objects and scenes that humans are also sensitive to, such as the depth and structure of surfaces (Li et al., 2023; Ranftl et al., 2022; Saxena et al., 2023; Liu et al., 2023; Goodwin et al., 2022; Tang et al., 2023; Amir et al., 2021; Chen et al., 2023; Bhattad et al., 2023; El Banani et al., 2024). Are the emergent capabilities of DNNs for 3D vision sufficient for solving VPT tasks?
@@
Here, we introduce the 3D perception challenge (3D-PC) to address this question and systematically compare 3D perceptual capabilities of humans and DNNs. The 3D-PC evaluates observers on (Fig 1): 1. identifying the order of two objects in depth (depth order), 2. predicting if one of two objects can “see” the other (VPT-basic), and 3. another version of VPT that limits the effectiveness of “shortcut” solutions (Geirhos et al., 2020a) (VPT-Strategy). The 3D-PC is distinct from existing psychological paradigms for evaluating VPT (Piaget et al., 1956; Bukowski, 2018; Martin et al., 2019) and computer vision challenges for 3D perception (El Banani et al., 2024; Amir et al., 2021) in two ways. First, unlike small-scale psychology studies of VPT, the 3D-PC uses a novel “3D Gaussian Splatting” (Kerbl et al., 2023) approach which permits the generation of endless real-world stimuli. Second, unlike existing computer vision challenges, our approach for data generation means that the 3D-PC tests and counterbalances labels for multiple 3D tasks on the exact same images, which controls for potential confounds in analysis and interpretation. We expect that DNNs which rival humans on the 3D-PC will become ideal models for a variety of real-world applications where machines must anticipate human behavior in real-time, as well as for enriching our understanding of how brains work (Fig. 1A).
@@
We built the 3D-PC and used it to evaluate 3D perception for human participants and 327 different DNNs. The DNNs we tested represented each of today’s leading approaches, from Visual Transformers (ViT) (Dosovitskiy et al., 2021a) trained on ImageNet-21k (Ridnik et al., 2021) to ChatGPT4 (Achiam et al., 2023) and Stable Diffusion 2.0 (Rombach et al., 2021).
@@
  • Fine-tuning the zoo of DNNs on VPT-basic boosted their performance to near human level. However, the performance of the DNNs — but not humans — dropped back to chance on VPT-Strategy. DNNs overly rely on the size and location features of objects instead of estimating their line-of-sight to solve VPT like humans do.
  • Our findings demonstrate that the visual strategies necessary for solving VPT do not emerge in DNNs from large-scale static image training or after directly fine-tuning on the task.We release the 3D-PC data, 3D Gaussian Splatting models, code, and human psychophysics to support the development of models that can perceive and reason about the 3D world like humans.
@@ 2 Related Work
The visual perception of 3D properties is a fundamentally ill-posed problem (Todd, 2004; Pizlo, 2010), which forces biological visual systems to rely on a variety of assumptions to decode the structure of objects and scenes. For example, variations in the lighting, texture gradients, retinal image disparity, and motion of an object all contribute to the perception of its 3D shape. 3D perception is further modulated by top-down beliefs about the structure of the world, which are either innate or shaped by prior sensory experiences, especially visual and haptic ones. In other words, humans learn about the 3D structure of the world in an embodied manner that is fundamentally different than how DNNs learn. In light of this difference, it would be remarkable if DNNs could accurately model how humans perceive their 3D world.
@@
VPT was devised to understand how capabilities for reasoning about objects in the world develop throughout the course of one’s life. At least two versions of VPT have been introduced over the years (Michelon & Zacks, 2006; Pizlo, 2022). The version of VPT that we study here — known in the developmental literature as “VPT-1” — is the more basic form, which is thought to rely on automatic feedforward processing in the visual system (Michelon & Zacks, 2006). In light of the well-documented similarities between feedforward processing in humans and DNNs (Serre, 2019; Kreiman & Serre, 2020), we reasoned that this version of VPT would maximize the chances of success for today’s DNNs.
3D perception in DNNs.
As deep neural networks (DNNs) have increased in scale and training dataset size over the past decade, their performance on essentially all visual challenges has improved. For example, there has also been significant progress in training DNNs to solve 3D computer vision tasks. For example, it is possible to train models that can precisely decode object and scene depth Yang et al. (2024), generate a high-resolution 3D model of an environment Kerbl et al. (2023), or even predict new camera views of an object given a single video of it Zhang et al. (2024). DNNs represent the state-of-the-art approach to nearly every 3D vision task today.
Surprisingly, there is also growing evidence that 3D perception emerges in DNNs even when they are trained on static image tasks. For example, DNNs trained with a variety of self-supervised learning techniques on static image datasets learn to represent the depth, surface normals, and 3D correspondence of features in scenes (Ranftl et al., 2022; Saxena et al., 2023; Liu et al., 2023; Goodwin et al., 2022; Tang et al., 2023; Amir et al., 2021; Chen et al., 2023; Bhattad et al., 2023; El Banani et al., 2024). While similarities between DNNs and human 3D perception have yet to be evaluated systematically, it has been shown that there are differences in how the two reason about the 3D shape of objects (Qian & Ullman, 2024). The 3D-PC complements prior work by systematically evaluating which aspects of human 3D perception today’s DNNs can and cannot accurately represent.
@@
Over recent years, DNNs have grown progressively more accurate as models of human vision for object recognition tasks (Geirhos et al., 2021; 2020a). At the same time, these models which succeed as models of human object recognition struggle to capture other aspects of visual perception (Bowers et al., 2022) including contextual illusions (Linsley et al., 2020), perceptual grouping (Kim* et al., 2020; Linsley et al., 2021), and categorical prototypes (Golan et al., 2020). There are also multiple reports showing that DNNs are growing less aligned with the visual strategies of humans and non-human primates as they improve on computer vision benchmarks (Linsley et al., 2023b; Fel* et al., 2022; Linsley et al., 2023a). The 3D-PC provides another axis upon which the field can evaluate DNNs as models of human vision.
@@ 3 Methods
To enable a fair comparison between human observers’ and DNNs’ 3D perceptual capabilities, we designed the 3D-PC framework with two goals: 1. posing different 3D tasks on the same set of stimuli, and 2. generating a large number of stimuli to properly train DNNs on these tasks. We achieved these goals by combining 3D Gaussian Splatting (Kerbl et al., 2023), videos from the Common Objects in 3D (Co3D) (Reizenstein et al., 2021) dataset, and Unity (Juliani et al., 2018; Pranckevicius, 2023) into a flexible data-generating framework.
@@
Our approach makes it possible to generate an unlimited number of visual stimuli that test an observer’s ability to solve complementary 3D perception tasks (depth order and VPT) while keeping visual statistics constant and ground truth labels counterbalanced across tasks. For the version of 3D-PC used in our evaluation and released at https://huggingface.co/datasets/3D-PC/3D-PC, the depth order and VPT-basic tasks are posed on the same set of 7,480 training images of 20 objects and scenes, and a set of 94 test images of 10 separate objects and scenes (Fig. 2). We held out a randomly selected 10% of the training images for validation and model checkpoint selection.
To build the VPT-Strategy task, we rendered images where we fixed the scene camera while we moved the green camera and red ball objects to precisely change the line-of-sight between them from unobstructed to obstructed and back. We reasoned that this experiment would reveal if an observer adopts the visual strategy of taking the perspective of the green camera, which is thought to be used by humans (Michelon & Zacks, 2006), from other strategies that relied on less robust feature-based shortcuts. This dataset consisted of a test set of 100 images for 10 objects and scenes that were not included in depth order or VPT-basic.
@@
Participants were shown instructions for one of the 3D-PC tasks, then provided 20 training examples to ensure that they properly understood it (Appendix Fig 8). These human training examples were drawn from the DNN training set. Each experimental trial consisted of the following sequence of events overlaid onto a white background: 1. a fixation cross displayed for $1000ms$; 2. an image displayed for $3000ms$, during which time the participants were asked to render a decision. Participants pressed one of the left or right arrow keys on their keyboards to provide decisions.
@@
We evaluated a wide range of DNNs on the 3D-PC, which represented the leading approaches for object classification, self-supervised pretraining, image generation, depth prediction, and vision language modeling (VLM). Our zoo includes 317 DNNs from PyTorch Image Models (TIMM) (Wightman, 2019), ranging from classic models like AlexNet (Krizhevsky et al., 2012a) to state-of-the-art models like EVA-02 (Fang et al., 2023) (see Appendix 1 for the complete list). We added foundational vision models like MAE (He et al., 2022), DINO v2 (Oquab et al., 2023), iBOT (Zhou et al., 2021), SAM (Kirillov et al., 2023), and Midas (Ranftl et al., 2022) (obtained from the GitHub repo of El Banani et al. (2024)). We also included Depth Anything (Yang et al., 2024), a foundational model 3D scene analysis and depth prediction (El Banani et al., 2024), as well as the Stable Diffusion 2.0 (Rombach et al., 2021) image generation model. Finally, we added state-of-the-art large vision language models (VLMs) ChatGPT4 (Achiam et al., 2023), Gemini (Team et al., 2023), and Claude 3 (Anthropic, 2024). We evaluated a total of 327 models on the 3D-PC.
@@
We evaluated all models except for the VLMs on the depth order and VPT-basic tasks in this challenge by training linear probes on image embeddings from their penultimate layers. Linear probes were trained using PyTorch (Paszke et al., 2019) for 50 epochs, a $5e$-4 learning rate, and early stopping (see Appendix A.5 for details). Training took approximately 20 minutes per model using NVIDIA-RTX 3090s. We tested the Stable Diffusion 2.0 model by adopting the evaluation method used in Li et al. (2023) (see Appendix A.8 for details). We evaluated the VLMs by providing them the same instructions and training images (along with ground truth labels) given to humans, then recording their responses to images from each task via model APIs.
@@ 4 Results
All humans participants and models were test on the same 94 images, each of which had a VPT label and a depth label. Human participants were on average 74.73% accurate at determining the depth order of objects, and 86.82% accurate at solving the VPT-basic task (Fig. 3; $p<0.001$ for both; statistical testing done through randomization tests (Edgington, 1964)). Humans were also significantly more accurate at solving VPT-basic than they were at the depth order task.
@@
DNNs showed the opposite pattern of results on depth order and VPT-basic tasks as humans after linear probing or prompting (Fig. 4): 15 of the DNNs we tested fell within the human accuracy confidence interval on the depth order task, and three even outperformed humans (Fig. 4A). In contrast, while humans were on average 86.82% accurate at VPT-basic, the DNN which performed the best on this task, the ImageNet 21K-trained beit (Bao et al., 2021a), was 53.82% accurate. Even commercial VLMs struggled on VPT-basic and were around chance accuracy (ChatGPT4: 52%, Gemini: 52%, and Claude 3: 50%). The depth order task was significantly easier for DNNs than VPT-basic ($p<0.001$), which is the opposite of humans (Fig. 4B).
@@
What drives the development of 3D perception in DNNs trained on static images? We hypothesized that as DNNs scale up, they learn ancillary strategies for processing natural images, including the ability to analyze the 3D structure of scenes. To investigate this possibility, we focused on the TIMM models in our DNN zoo. These models have previously been evaluated for object classification accuracy on ImageNet, which we used as a stand-in for DNN scale (Fel* et al., 2022; Linsley et al., 2023a; b). Consistent with our hypothesis, we found a strong and significant correlation between DNN performance on ImageNet and depth order task accuracy ($\rho=0.66,p<0.001$, Fig. 4C). Despite the very low accuracy of DNNs on VPT-basic, there was also a weaker but still significant correlation between performance on this task and ImageNet ($\rho=0.34,p<0.001$, the difference in correlations between the tasks is $\rho=0.32,p<0.001$; Fig. 4D). These results suggest that monocular depth cues develop in DNNs alongside their capabilities for object classification [^1] More work is needed to identify a causal relationship between the development of monocular depth cues and object recognition accuracy.. However, the depth cues that DNNs learn are poorly suited for VPT.
Figure 6: Even DNNs fine-tuned on VPT-basic fail on VPT-Strategy. (A) To better characterize the strategy used by humans and DNNs to solve VPT, we devised a new test, VPT-Strategy, in which the green camera and red ball are moved through a scene while holding the scene camera and a centrally-positioned object still. This task is easily solvable if an observer estimates the line-of-sight of the green camera; other strategies, such as those that rely on specific image features like the relative sizes and positions of objects (feature based), may be less effective. (B) Examples of VPT-Strategy stimuli along with the ground-truth label (top-row) and predictions by a ViT large after linearly probing or fine-tuning for VPT-basic (bottom-row). Decision attribution maps from each version of the ViT large, derived from “smooth gradients” (Smilkov et al., 2017), are overlaid onto bottom-row images (purple/blue=linearly probed, yellow/green=fine-tuned). The fine-tuned ViT locates the green camera and red ball but renders incorrect decisions. (C) DNNs fine-tuned on VPT-basic fail to solve VPT-Strategy; they rely on a brittle feature-based strategy. Humans, on the other hand, are 87% accurate; they likely estimate line-of-sight.
@@
One possible explanation for the failure of today’s DNNs on VPT-basic is that the task requires additional cues for 3D vision that cannot be easily learned from static images. To explore this possibility, we fine-tuned each of the TIMM models in our DNN zoo to solve depth order and VPT-basic (Fig. 5A). There was still a significant difference between DNN performance on the two tasks (Fig. 5B, $p<0.001$), but fine-tuning caused 97% of the DNNs to exceed human accuracy on depth order, and four of the DNNs to reach human accuracy on VPT-basic. DNN performance on the tasks more strongly correlated with ImageNet accuracy after fine-tuning than linear probing (compare Fig. 5C/D and Fig. 4C/D). We also compared the errors these DNNs made on both tasks to humans. We found nearly all of the fine-tuned DNNs were aligned with humans on depth order, and a handful were aligned with humans on VPT-basic (Fig. 9).
@@
Prior work has found that Deep Neural Networks (DNNs) can achieve human-level performance on visual tasks while using fundamentally different strategies than humans (Linsley et al., 2023b; Fel* et al., 2022; Linsley et al., 2023a). To investigate whether this issue applies to VPT, we designed a new experiment which we call VPT-Strategy.
VPT-Strategy was inspired by past work in developmental psychology, where it has been proposed that humans solve VPT by estimating line-of-sight (Fig. 6A) because humans respond in predictable ways when objects in a scene are slightly repositioned (Pizlo, 2022; Michelon & Zacks, 2006). In VPT-Strategy, observers solve the VPT task on a series of images rendered from a fixed camera viewpoint which show the green camera and red ball incrementally repositioned from one side of the screen to the other, passing by an occluding object in the process. This makes it possible to precisely map out the moments at which the green camera’s view of the red ball is unoccluded, then occluded, then unoccluded once more. DNNs that have been fine-tuned on VPT-basic behaved differently than humans on this task: humans were 87% accurate, but the highest performing DNN, the Swin Transformer (Liu et al., 2021) trained on ImageNet-21k, was only 66% accurate (Fig. 6C). Since DNNs cannot generalize to VPT-Strategy, it means that they do not learn to estimate line-of-sight to solve VPT.
To better understand the source of DNN failures, we derived “smooth gradients” (Smilkov et al., 2017) decision attribution maps of every DNN that we tested on VPT-Strategy, and then measured the overlap of these maps with the green camera and red ball objects using Dice similarity coefficient (Dice, 1945). These attribution maps revealed that the DNNs we evaluated did indeed learn to attend to the locations of the green camera and red ball objects when trying to solve VPT (Fig. 6B and Fig. 10), and the precision of their localization increased as a function of accuracy on VPT-basic (Fig. 10). When considered alongside the poor DNN performance on VPT-strategy, these results imply that DNNs learn a brittle feature-based strategy that is dependent on object properties like size and location instead of estimating the line-of-sight of the green camera like humans do (Fig. 6C).